Calculate the mass of solid E inside the cone z =sqr( ?4x^2 + 4y^2 ) in the
sphere of radius 4 whose density is given by δ(x,y,z)=16−x^2 −y^2 −z^2
Answer #1

2 Bol: cone: Z = V 4 GByzj o .. 100 opher x² + y² 72²=42. - 56419, 2) = 16-(x?ty?rz) intersection of sphere and cone - - - -so schy, z) = 16-82 x=8 sino caso 9 = Phiлф 4 Z=P cood 033 34 050 5 2 035 o sind(345) 4 Mass = ss savai l s sino dedo de sint